9.Consider the following gaseous equilibrium:
Kp = 2.59 at 128 oC. What is G for the reaction?
10. A 0.55 g of sodium oxalic acid (NaHC2O4, pKa = 4.19) and 0.65 g of disodium oxalate (Na2C2O4) were dissolved in 125 mL. What is the pH of the solution?
11. 3-(Cyclohexylamino)-2-hydroxy-1-propanesulfonic acid is the name of one of a Goodes buffers, which is why it is given the acronym: CAPSO, a conjugate acid. The molecular weight of CAPSO is 237.32. The pKa of CAPSO is 10.4. How would you prepare 800 mL of 0.125 M CAPSO buffer pH 10.2? There is more than one answer to this question
Answer #1

9) The free energy change is calculated:

ΔG = - R * T * ln K = - 0.008314 * 401 K * ln (2.59) = - 3.18 kJ / mol

10) The moles of the buffer components are calculated:

n Acid = g / MM = 0.55 / 112 = 0.005 mol

n Salt = 0.65 / 134 = 0.005 mol

The pH = pKa = 4.19

11) The molar ratio of the buffer components is calculated:

n Salt / n Acid = 10 ^ (pH - pKa) = 10 ^ (10.2 - 10.4) = 0.6

It has:

i) n Salt - 0.6 * n Acid = 0

ii) n Salt + n Acid = M * V = 0.125 M * 0.8 L = 0.1 mol

The system of equations is solved:

n Salt = 0.04 mol

n Acid = 0.06 mol

A CAPSO mass of:

m CAPSO = n * MM = 0.1 mol * 237.32 g / mol = 2.37 g

0.1 M NaOH is added a volume of:

V NaOH = n * 1000 / M = 0.04 * 1000 / 0.1 = 400 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.